OpenStudy (anonymous):

why does |-3b|=9 and |3b|=9 is the same answer? like how do u solve it ?

OpenStudy (anonymous):

Mod of any number is its positive value. For example, mod of 4 is 4 and mod of -4 is 4

OpenStudy (anonymous):

thanks so much for ur help :D

OpenStudy (anonymous):

You're welcome :) You can solve it now?

OpenStudy (anonymous):

I understand why but I don't know how to solve it .. lol

OpenStudy (anonymous):

Oh okay mod of 3b = 9 b = 9/3 = 3 OR and mod of -3b = 9 b = 9/(-3) = -3 So b = 3 or -3

OpenStudy (anonymous):

Tell me if you get it, or I'll try to make it clearer...

OpenStudy (anonymous):

@miss.pinkcandyysXO

OpenStudy (anonymous):

yea I got it . u divide 3 to 9 and -3 to 9 right ?

OpenStudy (anonymous):

Right

OpenStudy (anonymous):

thanks :D

OpenStudy (anonymous):

You're most welcome! :D